OpenStudy (mrbaconqwerty):

WILL MEDAL!!! What is this question asking me to do. I'm really confused: A room with width w, length L, and height h with 4 walls needs to be painted. Suppose L is equal to w. Write a formula for A in terms of w and h.

OpenStudy (mikey):

does it give you any numbers at all

OpenStudy (croit):

I assume they want a general formula. What does the A stand for?

OpenStudy (mikey):

I think A is the whole room

OpenStudy (croit):

The volume of the room or just the area of all the walls?

OpenStudy (mikey):

good question but idk

jhonyy9 (jhonyy9):

A mean the total area what need be painting so these 4 walls and the top area hope is understandably now

jhonyy9 (jhonyy9):

so than L=w and is needed painting just this 4 walls so 1 wall has area = L*w but bc. w=L so will be 1 wall area = L*L = L^2

jhonyy9 (jhonyy9):

so than you know area of 1 wall just you need multiplie this by 4 and will get the right equation yes ?

jhonyy9 (jhonyy9):

ohhh sorry now i see that L = w but you need writing this equation using w and h

jhonyy9 (jhonyy9):

so area in term of w and h will be area 1 wall = w*h

OpenStudy (mrbaconqwerty):

No worries! I just appreciate you trying to help :) @jhonyy9

jhonyy9 (jhonyy9):

ok. ty .- but this above wrote is right

OpenStudy (mrbaconqwerty):

Ok. Thank you so much! :) @jhonyy9

jhonyy9 (jhonyy9):

do you understand it now sure ? than how will be for these 4 walls

jhonyy9 (jhonyy9):

hope you understand it - just multiplie this equation by 4

OpenStudy (mrbaconqwerty):

Oh! So then it would be 4wh? Because there are 4 walls?

jhonyy9 (jhonyy9):

exactly very nice - so this mean that you understood it now sure right good

OpenStudy (mrbaconqwerty):

Alright, thanks!

jhonyy9 (jhonyy9):

np welcome and anytime good luck bye bye